# SADA: Semantic adversarial unsupervised domain adaptation for Temporal Action Localization

### v2

@@ 3.6 Training
Note again is optimized with domain while promotes the alignment between both domains and . Moreover, and are tunable parameters that control their influence on the loss.
@@ 4.1 Benchmarks for UDA on sparse TAL
Evaluating domain adaptation-based methods in the context of video understanding is a challenging issue that requires the definition of a reasonable domain gap and identifying a sufficiently large set of intersecting action classes. Dividing existing datasets into different domains that comply with these conditions often restricts the amount of data to learn and adapt. SSTDA [7] approaches this problem on GTEA [8] and Breakfast [9] by defining a subject-based partitioning where they aim to adapt the model to new unseen subjects. However, as little data is available from a single subject in those datasets, they group several users for training. This allows the model to generalize over the subject variability under study, making it unsuitable to test for domain adaptation. Closely related to our work, [61, munro2020multi] propose several UDA scenarios for video classification based on EpicKitchens100 [14]. These define 3 domains based on the data of 3 different kitchens, thus performing cross-kitchen evaluation. This limits even further the amount of data in each domain –i.e., between 15 to 29 videos per domain.
EpicKitchens100: To overcome these limitations, we first propose a new set of 6 different scenarios (S1, …, S6) for sparse TAL based on the EpicKitchen100 [14] (see Fig. 3). EpicKitchen100 presents an ideal base for our tasks as it has become a gold standard to evaluate complex sparse detection scenarios on long egocentric videos (up to 45 minutes). We identify two domain gaps in this dataset: an appearance domain shift based on the different colors of the kitchen counters; and an acquisition domain shift that results in the differences of lighting and camera conditions when extending EpicKitchens55 [14] into its new version EpicKitchens100 [14]. This permits to define a rich set of benchmarks that provide a more fine-grained and comprehensive evaluation. Importantly, this strategy is also suitable for single-source settings which do not allow an easy generalization over the shift under study. For example, if a model is trained with dark-counter kitchens only, it cannot easily generalize to white-counter kitchens.
@@
One important consideration is that EpicKitchens100 exhibits a very strong long-tail label distribution, where the vast majority of its 97 actions represent only a marginal percentage of the overall data (see Supp.). Adapting DA methods to these long tail distributions falls beyond the scope of this paper and most of the existing literature. For this reason, we follow a similar approach to that proposed by [munro2020multi] and limit our evaluation to the 10 majority classes representing of the original dataset. This still results in a very complex task proof of which is that the best-performing models to date attain less than a 30% mAP [21, 10].
CharadesEgo: One important aspect of the 6 previous setups is that all the domains share an egocentric viewpoint. For this reason, we evaluate the model degradation under extreme domain shifts caused by major changes in the perspective. For this, we use CharadesEgo [13] which extends the original third-person videos from Charades [62] by matching them with their corresponding egocentric videos. Following [12], we define the Source domain as the third-person videos, and the Target domain as the egocentric ones. Given the extreme shift that these present (see Fig. 4), we limit the task to predicting the verb of each of the action segments, and similarly to the previous case, we keep the 10 majority verbs only. This mitigates the effect of long-tail action distributions in our evaluation which falls beyond the scope of this paper. See the Supp. for more details.
@@ 4.2 Experimental results
| Scenario | Model | mAP10% | mAP20% | mAP30% | mAP40% | mAP50% | Avg |
[…]
| Tridet [10] | 29.47 | 28.32 | 25.50 | 21.99 | 16.34 | 24.32 | |
| Ours (source-only) | 30.03 | 28.70 | 26.62 | 23.03 | 17.79 | 25.23 | |
| Actionformer+SADA [21] | 30.54 | 29.31 | 27.36 | 23.75 | 18.80 | 25.95 | |
| Tridet+SADA [10] | 31.34 | 30.16 | 27.94 | 24.48 | 19.31 | 26.64 | |
| Ours (SADA) | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 | |
[…]
| Tridet [10] | 29.87 | 28.39 | 25.97 | 22.06 | 16.94 | 24.65 | |
| Ours (source-only) | 30.74 | 29.47 | 27.23 | 23.53 | 18.07 | 25.80 | |
| Actionformer+SADA [21] | 31.71 | 30.30 | 27.95 | 24.08 | 18.69 | 26.55 | |
| Tridet+SADA [10] | 29.55 | 28.27 | 26.16 | 22.89 | 17.63 | 24.90 | |
| Ours (SADA) | 31.60 | 30.29 | 28.22 | 24.47 | 18.98 | 26.72 | |
[…]
| Tridet [10] | 34.01 | 32.52 | 30.07 | 26.40 | 19.78 | 28.55 | |
| Ours (source-only) | 34.41 | 33.38 | 30.58 | 26.99 | 21.00 | 29.27 | |
| Actionformer+SADA [21] | 34.11 | 32.87 | 30.60 | 27.05 | 20.93 | 29.11 | |
| Tridet+SADA [10] | 34.50 | 33.19 | 30.65 | 27.25 | 20.83 | 29.29 | |
| Ours (SADA) | 34.86 | 33.73 | 31.16 | 27.45 | 21.46 | 29.73 | |
[…]
| Tridet [10] | 30.03 | 28.97 | 26.96 | 23.48 | 18.18 | 25.52 | |
| Ours (source-only) | 29.65 | 28.69 | 26.86 | 23.88 | 19.14 | 25.64 | |
| Actionformer+SADA [21] | 29.84 | 28.89 | 26.83 | 23.76 | 18.85 | 25.63 | |
| Tridet+SADA [10] | 30.21 | 29.33 | 27.58 | 24.35 | 19.44 | 26.18 | |
| Ours (SADA) | 31.54 | 30.68 | 28.77 | 25.52 | 20.22 | 27.34 | |
[…]
| Tridet [10] | 24.77 | 22.93 | 21.49 | 19.09 | 15.15 | 20.48 | |
| Ours (source-only) | 25.58 | 24.79 | 23.08 | 19.56 | 15.15 | 21.63 | |
| Actionformer+SADA [21] | 24.85 | 23.99 | 22.21 | 19.11 | 15.28 | 21.09 | |
| Tridet+SADA [10] | 24.88 | 24.00 | 22.20 | 19.02 | 14.88 | 21.00 | |
| Ours (SADA) | 25.93 | 25.06 | 23.47 | 20.45 | 16.12 | 22.21 | |
[…]
| Tridet [10] | 22.47 | 21.57 | 20.19 | 17.87 | 14.41 | 19.30 | |
| Ours (source-only) | 20.96 | 20.22 | 19.08 | 16.97 | 14.09 | 18.27 | |
| Actionformer+SADA [21] | 23.05 | 22.10 | 20.71 | 18.31 | 14.72 | 19.78 | |
| Tridet+SADA [10] | 21.00 | 20.08 | 18.64 | 16.39 | 13.13 | 17.85 | |
| Ours (SADA) | 23.94 | 22.95 | 21.47 | 19.16 | 15.24 | 20.55 |
@@
EpicKitchens100. In Table I we first present the results obtained in the 4 different scenarios that we designed to evaluate the performance of our method when facing different appearance shifts induced by changes in the background information. Concretely, we first evaluate the performance of the Actionformer [21] and TriDet [10], the two best-performing methods on EpicKitchens100 dataset [14], as well as our proposed architecture without our SADA loss. We then compare these 3 architectures with their extensions which include the SADA loss. Observe that our proposed loss improves the source-only version of all the architectures in all these 4 scenarios, improving by up to a mAP when evaluating our architecture on S1. Additionally, we observe that our final architecture –i.e., Ours(SADA)– yields an improvement over the second-best performing source-only method of up to a mAP for the black-counter kitchens scenarios and of mAP for the white-counter kitchens ones.
Similarly, the last two scenarios of Table I depict the setups containing acquisition domain shifts. These scenarios present a similar behavior than before. In all but one case on the Tridet [10] architecture, the use of the SADA loss yields a performance gain of up to mAP. This is a relative improvement. Our final model, moreover, improves the second-best performing source-only baselines by mAP and mAP, respectively.
| Model | mAP10% | mAP20% | mAP30% | mAP40% | mAP50% | Avg |
@@
| Ours (source-only) | 30.68 | 27.74 | 23.41 | 18.46 | 13.58 | 22.77 |
| Actionformer+SADA [21] | 31.46 | 28.57 | 24.17 | 19.04 | 13.87 | 23.42 |
| Tridet+SADA [10] | 31.53 | 28.41 | 24.04 | 18.88 | 13.98 | 23.37 |
@@
CharadesEgo. In Table II we report a similar experimental comparison of our proposed loss when evaluated on CharadesEgo. Concretely, we showcase the performance boost that SADA reports on the three test architectures –i.e., Actionformer [21], Tridet [10] and Ours. In the case of Tridet [10], for instance, the improvement reaches a mAP. In our proposed architecture, SADA yields an improvement over the source-only version of mAP. We additionally highlight that our proposed architecture Ours (SADA) attains the state-of-the-art results over all the baseline methods.
@@ 4.3 Ablation studies
Comparing to other domain adaptation methods. In Sec. 4.2 we showed that SADA consistently improves the performance of the three tested state-of-the-art source-only architectures when evaluated on our newly proposed setups. The question remains however of how well does SADA perform compared to existing domain adaptation methods. In this regard, following existing video-based domain adaptation works, we first evaluate various canonical UDA domain adaptation methods –i.e., ADDA [40], WDGRL [42], DANN [1] and MSTN [2]. These methods are integrated into our proposed underlying architecture resulting in a fair comparison with SADA. Additionally, we find that to the best of our knowledge, there is no existing UDA method for TAL in the literature that is directly comparable to this work. Nevertheless, to provide a richer comparison, we adapt the closest action segmentation proposal, SSTDA [7] and a state-of-the-art domain-adaptation method for video classification –i.e., TranSVAE [61]– to our proposed setup. We refer to the Supp. for all the details of these baselines.
Concretely, in Table III we compare SADA with these baselines on S1 and S2. These results empirically demonstrate the effectiveness of our method by attaining the best results over all tested UDA methods. More in detail, SADA yields an improvement of up to mAP and mAP on each of the scenarios respectively, and a mAP and mAP over the second best-performing method.
[…]
@@
| Model | mAP {10,30,50}% | Avg | ||
| ADDA [40] | 30.55 | 27.09 | 18.17 | 25.27 |
| WDGRL [42] | 24.85 | 21.87 | 13.68 | 20.13 |
| DANN [1] | 30.87 | 27.45 | 18.29 | 25.97 |
| MSTN [2] | 30.88 | 27.44 | 18.25 | 25.52 |
| SSTDA [7] | 29.92 | 26.49 | 17.55 | 24.65 |
| TranSVAE [65] | 25.88 | 21.74 | 13.12 | 20.25 |
| Ours (SADA) | 31.60 | 28.22 | 18.98 | 26.27 |
Analysis of our loss. Next, we ablate over different variants of our proposed SADA loss (see Eq. 9). Concretely, in Table IV we study the effect of class-wise distribution alignment (Eq. 7), global distribution alignment (equivalent to DANN [1]) and semantic background alignment (Eq. 8). In this regard, we highlight that the global adaptation consistently improves upon aligning only background embeddings. This is because the latter yields only a partial alignment of the embeddings, not considering class anchors. Therefore, a rougher yet complete adaptation might seem beneficial. We also observe that aligning local class-wise distributions has a considerable positive effect, especially in the first scenario. Its effect, however, considerably decreases when combined with a global alignment [1], as all the class anchors are then subject to the concurrent alignments of domain-level and class-wise distributions. This observation is also consistent with the performance decrease that we observe when combining global and background alignment, which again suggests that the concurrent alignment of background embeddings with two adaptation losses is harmful to performance. Finally, we observe that we consistently obtain the best results when the local alignment loss is coupled with the complementary (non-overlapping) background loss – i.e., SADA loss – indicating that this semantic fine-grained, yet complete, alignment is the most desirable approach.

### v3

@@ 3.6 Training
Note again is optimized with domain while promotes the alignment between both domains and . Moreover, and are tunable parameters.
@@ 4.1 Benchmarks for UDA on sparse TAL
Evaluating domain adaptation-based methods in the context of video understanding is a challenging issue that requires the definition of a reasonable domain gap and identifying a sufficiently large set of intersecting action classes. Dividing existing datasets into different domains that comply with these conditions often restricts the amount of data to learn and adapt. SSTDA [8] approaches this problem on GTEA [16] and Breakfast [29] by defining a subject-based partitioning where they aim to adapt the model to new unseen subjects. However, as little data is available from a single subject in those datasets, they group several users for training. This allows the model to generalize over the subject variability under study, making it unsuitable to test for domain adaptation. Closely related to our work, [57, 41] propose several UDA scenarios for video classification based on EpicKitchens100 [14]. These define 3 domains based on the data of 3 different kitchens, thus performing cross-kitchen evaluation. This limits even further the amount of data in each domain – i.e., between 15 to 29 videos per domain.
EpicKitchens100: To overcome these limitations, we first propose a new set of 6 different scenarios (S1, …, S6) for sparse TAL based on EK100 [14], (see Fig. 3). EK100 presents an ideal base for our tasks as it has become a gold standard to evaluate complex sparse detection scenarios on long egocentric videos (up to 45 minutes). We identify two domain gaps in this dataset: an appearance domain shift based on the different colors of the kitchen counters; and an acquisition domain shift that results in the differences of lighting and camera conditions when extending EK55 [14] into its new version EK100 [14]. This permits to define a rich set of benchmarks that provide a more fine-grained and comprehensive evaluation. Importantly, this strategy is also suitable for single-source settings which do not allow an easy generalization over the shift under study. For example, if a model is trained with dark-counter kitchens only, it cannot easily generalize to white-counter kitchens.
@@
One important consideration is that EK100 exhibits a very strong long-tail label distribution, where the vast majority of its 97 actions represent only a marginal percentage of the overall data (see Supp.). Adapting DA methods to these long tail distributions falls beyond the scope of this paper and most of the existing literature. For this reason, we follow a similar approach to [41] and limit our evaluation to the 10 majority classes for both domains and , representing of the original data. This still results in a very complex task, proof of which is that the best-performing models to date attain less than a 30% mAP [63, 48].
CharadesEgo: One important aspect of the 6 previous setups is that all the domains share an egocentric viewpoint. For this reason, we evaluate the model degradation under extreme domain shifts caused by major changes in the perspective. For this, we use CharadesEgo [50] which extends the original third-person videos from Charades [51] by matching them with their corresponding egocentric videos. Following [64], we define the Source domain as the third-person videos, and the Target domain as the egocentric ones. Given the extreme shift that these present, we limit the task to predicting the verb of each of the action segments, and similarly to the previous case, we keep the 10 majority verbs only. We refer to the Supp. for more details.
@@ 4.2 Experimental results
Scenario | Model | mAP {10,20,30,40,50}% | Avg | Scenario | Model | mAP {10,20,30,40,50}% | Avg | ||||||||
S1 | Actionformer [63] | 30.21 | 28.73 | 26.39 | 22.60 | 17.09 | 25.00 | S2 | Actionformer [63] | 27.46 | 26.54 | 24.61 | 21.85 | 17.19 | 23.53 |
Tridet [48] | 29.87 | 28.39 | 25.97 | 22.06 | 16.94 | 24.65 | Tridet [48] | 30.03 | 28.97 | 26.96 | 23.48 | 18.18 | 25.52 | ||
Ours (src-only) | 30.74 | 29.47 | 27.23 | 23.53 | 18.07 | 25.80 | Ours (src-only) | 29.65 | 28.69 | 26.86 | 23.88 | 19.14 | 25.64 | ||
Actionformer+SADA | 31.71 | 30.30 | 27.95 | 24.08 | 18.69 | 26.55 | Actionformer+SADA | 29.84 | 28.89 | 26.83 | 23.76 | 18.85 | 25.63 | ||
Tridet+SADA | 29.55 | 28.27 | 26.16 | 22.89 | 17.63 | 24.90 | Tridet+SADA | 30.21 | 29.33 | 27.58 | 24.35 | 19.44 | 26.18 | ||
Ours (SADA) | 31.60 | 30.29 | 28.22 | 24.47 | 18.98 | 26.72 | Ours (SADA) | 31.54 | 30.68 | 28.77 | 25.52 | 20.22 | 27.34 | ||
S3 | Actionformer [63] | 28.11 | 26.94 | 24.89 | 21.43 | 16.51 | 23.57 | S4 | Actionformer [63] | 33.52 | 32.31 | 29.84 | 26.48 | 20.11 | 28.45 |
Tridet [48] | 29.47 | 28.32 | 25.50 | 21.99 | 16.34 | 24.32 | Tridet [48] | 34.01 | 32.52 | 30.07 | 26.40 | 19.78 | 28.55 | ||
Ours (src-only) | 30.03 | 28.70 | 26.62 | 23.03 | 17.79 | 25.23 | Ours (src-only) | 34.41 | 33.38 | 30.58 | 26.99 | 21.00 | 29.27 | ||
Actionformer+SADA | 30.54 | 29.31 | 27.36 | 23.75 | 18.80 | 25.95 | Actionformer+SADA | 34.11 | 32.87 | 30.60 | 27.05 | 20.93 | 29.11 | ||
Tridet+SADA | 31.34 | 30.16 | 27.94 | 24.48 | 19.31 | 26.64 | Tridet+SADA | 34.50 | 33.19 | 30.65 | 27.25 | 20.83 | 29.29 | ||
Ours (SADA) | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 | Ours (SADA) | 34.86 | 33.73 | 31.16 | 27.45 | 21.46 | 29.73 | ||
S5 | Actionformer [63] | 22.87 | 21.87 | 20.10 | 17.23 | 13.33 | 19.08 | S6 | Actionformer [63] | 22.16 | 21.22 | 19.71 | 17.44 | 14.08 | 18.92 |
Tridet [48] | 24.77 | 22.93 | 21.49 | 19.09 | 15.15 | 20.48 | Tridet [48] | 22.47 | 21.57 | 20.19 | 17.87 | 14.41 | 19.30 | ||
Ours (src-only) | 25.58 | 24.79 | 23.08 | 19.56 | 15.15 | 21.63 | Ours (src-only) | 20.96 | 20.22 | 19.08 | 16.97 | 14.09 | 18.27 | ||
Actionformer+SADA | 24.85 | 23.99 | 22.21 | 19.11 | 15.28 | 21.09 | Actionformer+SADA | 23.05 | 22.10 | 20.71 | 18.31 | 14.72 | 19.78 | ||
Tridet+SADA | 24.88 | 24.00 | 22.20 | 19.02 | 14.88 | 21.00 | Tridet+SADA | 21.00 | 20.08 | 18.64 | 16.39 | 13.13 | 17.85 | ||
Ours (SADA) | 25.93 | 25.06 | 23.47 | 20.45 | 16.12 | 22.21 | Ours (SADA) | 23.94 | 22.95 | 21.47 | 19.16 | 15.24 | 20.55 |
@@
EpicKitchens100. In Tab. 1 we first present the results obtained in the 4 different scenarios that we designed to evaluate the performance of our method when facing different appearance shifts induced by changes in the background information. Concretely, we first evaluate the performance of the Actionformer [63] and TriDet [48], the two best-performing methods on EK100 dataset [14], as well as our proposed architecture without our SADA loss. We then compare these 3 architectures with their extensions which include the SADA loss. Observe that our proposed loss improves the source-only (SO) version of all the architectures in all these 4 scenarios, showing robustness across the chosen underlying architecture. For instance on S3, our proposed loss yields an improvement of up to mAP over its respective SO version. Additionally, we observe that our final architecture – i.e., Ours(SADA) – improves the best-performing existing SO method – i.e., Tridet[48] – by up to a mAP for the black-counter kitchens scenarios – i.e., S1 and S3 – and mAP for the white-counter kitchens –i.e., S2 and S4. Similarly, the last two scenarios of Tab. 1 present a similar behavior than before. In all but one case on the Tridet [48] architecture, the use of the SADA loss yields a performance gain of up to mAP. This is a relative improvement. Our final model, moreover, improves the best-performing existing baseline – i.e., Tridet – by mAP and mAP, respectively.
Model | mAP {10,20,30,40,50}% | Avg | ||||
@@
Ours (src-only) | 30.68 | 27.74 | 23.41 | 18.46 | 13.58 | 22.77 |
Actionformer+SADA | 31.46 | 28.57 | 24.17 | 19.04 | 13.87 | 23.42 |
Tridet+SADA | 31.53 | 28.41 | 24.04 | 18.88 | 13.98 | 23.37 |
@@
CharadesEgo. In Tab. 2 we report a similar experimental comparison of our proposed loss when evaluated on CharadesEgo. Concretely, we showcase the performance boost that SADA reports on the three test architectures – i.e., Actionformer [63], Tridet [48] and Ours. Observe that SADA, for instance, improves the performance of Tridet [48] by mAP. Similarly, our proposed architecture Ours (SADA) improves the SO version by mAP and yields the overall highest scores.
@@ 4.3 Ablation studies
Comparing to other domain adaptation methods. In Sec. 4.2 we showed that SADA consistently improves the performance of the three tested state-of-the-art SO architectures when evaluated on our newly proposed setups. The question remains however of how well does SADA perform compared to existing domain adaptation methods. In this regard, following existing video-based domain adaptation works, we first evaluate various canonical UDA domain adaptation methods –i.e., DANN [19], ADDA [55], WDGRL [47], MSTN [59], FGDA [20] and DRDA [24]. These methods are integrated into our proposed underlying architecture resulting in a fair comparison with SADA. Additionally, we find that to the best of our knowledge, there is no existing UDA method for TAL in the literature that is directly comparable to us. Nevertheless, to provide a richer comparison, we adapt the closest action segmentation proposal, SSTDA [8] and a state-of-the-art domain-adaptation method for video classification – i.e., TranSVAE [57] – to our proposed setup (we refer to the Supp. for all the details of all these baselines). Concretely, in Tab. 3 we compare SADA with these baselines on S3. These results empirically demonstrate the effectiveness of our method by attaining the best results over all tested UDA methods. More in detail, SADA yields an improvement of up to mAP over the second best-performing method (ADDA). Find the complete ablation for the other scenarios in the Supp.
@@
Analysis of our loss. Next, we ablate over different variants of our proposed SADA loss (see Eq. 9). Concretely, in Tab. 4 we study the effect of class-wise distribution alignment (Eq. 7), global distribution alignment (equivalent to DANN [19]) and semantic background alignment (Eq. 8). In this regard, we highlight that the global adaptation seems to consistently improve upon aligning only background embeddings. This is because the latter yields only a partial alignment of the embeddings, not considering class anchors. Therefore, a rougher yet complete adaptation might seem beneficial. We also observe that aligning local class-wise distributions has a considerable positive effect. Its effect, however, considerably decreases when combined with a global alignment [19], as all the class anchors are then subject to the concurrent alignments of domain-level and class-wise distributions. This observation is also consistent with the performance decrease that we observe when combining global and background alignment, which again suggests that the concurrent alignment of background embeddings with two adaptation losses is harmful to performance. Finally, we observe that we consistently obtain the best results when the local alignment loss is coupled with the complementary (non-overlapping) background loss – i.e., SADA loss – indicating that this semantic fine-grained, yet complete, alignment is the most desirable approach. Find in the Supp. the complete ablations for additional scenarios and the class-wise analysis.
Local | Global | Bkg | mAP {10,30,50}% | Avg | ||
30.03 | 26.62 | 17.79 | 24.81 | |||
30.48 | 27.07 | 18.12 | 25.22 | |||
30.34 | 26.66 | 17.04 | 24.68 | |||
29.76 | 26.63 | 17.52 | 24.64 | |||
31.36 | 28.01 | 18.67 | 26.01 | |||
30.09 | 26.88 | 17.43 | 24.80 | |||
32.69 | 29.17 | 19.72 | 27.19 |
Analysis of the impact of background anchors. One critical aspect of anchor-based methods for TAL is the presence of numerous background anchors. We argue that the confusion that these embeddings induce is one of the main challenges for the transfer to a different domain. Concretely, we hypothesize this is because of their high intra-class variance, and their low inter-class variance with other action classes –i.e., they share aspects like appearance. Nevertheless, in this ablation we show that SADA partially mitigates this effect. For this, we design an experiment that artificially masks out all the background anchors during inference to elucidate the ideal performance if we could ignore entirely the confusion they generate –i.e., by being wrongly predicted as an action. Concretely, in Tab. 5 we compare our proposed methods with its SO version, and observe that masking out all the background anchors yields an overall improvement of and mAP, respectively. Hence, SADA reduces the impact of masking the background anchors by mAP, indicating that our fine-grained alignment permits a better knowledge transfer to the target domain, mitigating the negative effect of these anchors. Find in the Supp. the complete ablation.
Method | Mask bkg anchors | mAP {10,30,50}% | Avg | Perf. gap | ||
Ours(src-only) | 30.03 | 26.62 | 17.79 | 24.81 | - | |
@@
Ours(SADA) | 32.69 | 29.17 | 19.72 | 27.19 | - | |